OpenStudy (anonymous):

two circles have areas of 16pi and 25pi. find the ratio of their circumferences

ganeshie8 (ganeshie8):

ratio of areas = \(\dfrac{16\pi}{25\pi} = \dfrac{16}{25}\) therefore ratio of curcumferences = \(\dfrac{\sqrt{16}}{\sqrt{25}} = \dfrac{4}{5}\)

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

Yw, that works more generally. You may use it for any kind of lengths : If the "areas" of similar figures are in ratio \(a:b\), then the ratio of their "lengths" will be in ratio \(\sqrt{a}:\sqrt{b}\)

OpenStudy (anonymous):

so it works for area and circumferenceÉ

OpenStudy (anonymous):

1st cirlcle area=>pi r*r=16pi ,so r*r=16 and r=4 similarly ,2nd circle area=> pi*r*r=25pi so,r=5 circumference of 1st circle=>2 pi*r so,area =2*pi*4 similarly 2nd area of circle => 2*pi*5 dividing both of them we will get=> 2*pi*4/2*pi*5=4/5 so,ratio is 4/5 :)

OpenStudy (anonymous):

thanks)